OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares for the numbers 45,77,112

OpenStudy (anonymous):

not sure what this means, but \(45=7^2-2^2\)

OpenStudy (anonymous):

is this the kind of answer you are looking for?

OpenStudy (anonymous):

since \(a^2-b^2=(a+b)(a-b)\) you can start with \((a+b)(a-b)=77\) for example since 77 factors as \(11\times 7\) that tells you \[a+b=11\] \[a-b=7\] and then you can solve for \(a\) and \(b\) by adding the two equations to get \[2a=18\] and so \(a=9\) making \(b=2\) and thus \[77=(9+2)(9-2)=9^2-2^2\]